OpenStudy (anonymous):

What is the largest area that can be enclosed with 400 feet of fencing? What are the dimensons of the rectangle?

OpenStudy (paxpolaris):

what grade question is this?

OpenStudy (anonymous):

Precal

OpenStudy (paxpolaris):

is that all the info. Perimeter=400

OpenStudy (abdullah1995):

ahh

OpenStudy (abdullah1995):

this question is easy :)

OpenStudy (abdullah1995):

i think it's optimization if i am not wrong ?

OpenStudy (abdullah1995):

2x+y = 400

OpenStudy (abdullah1995):

x = short side ; y = long side of rectangle

OpenStudy (anonymous):

What I thought it was 20=L+W

OpenStudy (abdullah1995):

one long side of rectangle is not included because in question it says 3 dimensions

OpenStudy (abdullah1995):

where do you get the 20 from ?

OpenStudy (anonymous):

200 my bad

OpenStudy (abdullah1995):

where did you get the 200 from ? lol

OpenStudy (anonymous):

well if you take 400=2X+2y and divide it by two

OpenStudy (paxpolaris):

right !!! ... L+W=200 would be if all four sides are using fence

OpenStudy (abdullah1995):

oh ok. i read the question wrong. yes you can say 200 = l + w

OpenStudy (anonymous):

so where do I go from there

OpenStudy (abdullah1995):

first tell us are all 4 sides using fence ? or just 3 ?

OpenStudy (anonymous):

4

OpenStudy (abdullah1995):

ok then it's 200 = l + w

OpenStudy (anonymous):

Okay I just dont know what to do after that how would I go about solving it

OpenStudy (abdullah1995):

well you have to form another equation relating L and W

OpenStudy (anonymous):

So like a function

OpenStudy (abdullah1995):

well. if you recall the area function, it is LxW = Area

OpenStudy (anonymous):

Yeah

OpenStudy (abdullah1995):

so from your first perimeter equation, separate either L or W

OpenStudy (abdullah1995):

lets say we separate L ; therefore L = 200 - W

OpenStudy (anonymous):

okay

OpenStudy (abdullah1995):

so now going back to the area equation, we see we have LxW = A ; but we know L = 200-W

OpenStudy (anonymous):

okay

OpenStudy (abdullah1995):

so A = (200-W)(W)

OpenStudy (anonymous):

Okay

OpenStudy (abdullah1995):

expand it out, A = 200W - W^2

OpenStudy (anonymous):

thats what I got too

OpenStudy (abdullah1995):

now set the area = 0 and take derivative

OpenStudy (anonymous):

I dont know deriviates I am only in precal

OpenStudy (abdullah1995):

oh ...

OpenStudy (abdullah1995):

ok then

OpenStudy (paxpolaris):

 for max area the shape has to be square